OpenStudy (anonymous):

help please Y = -2 x - 6 covert to slope-intercept form best answer gets medal

OpenStudy (nikato):

It's already in slope intercept form

OpenStudy (anonymous):

can i get help one more thing

OpenStudy (anonymous):

Y = 3 – x in slope-intercept form

OpenStudy (nikato):

That's also already int hat form already

OpenStudy (anonymous):

Yeah its already in slop-intercept form.

OpenStudy (anonymous):

reallymy teacher said You answer has to be in slope-intercept form. Only provide the correct answer and the work you used to find it.

OpenStudy (anonymous):

well ok iguess

OpenStudy (triciaal):

Y = 3 – x I would change this one around y = -x + 3 If the teacher wants more put this for the first one the first one Y = -2x + (-6)

OpenStudy (triciaal):

the first one slope = -2 intercept = -6 the 2nd one m = -1 and intercept = 3

OpenStudy (triciaal):

what really was the complete original question?

OpenStudy (xapproachesinfinity):

that is just unnecessary! your equation both are in slope intercept form doesn't matter if you do this y=-x+3 or y=3-x these are all intercept form of a line